Hi,I need help on Finding Missing Numbers in Patterns and I was wondering do you know how they got from number to number in this problem:80,77,71, blank,50,35,17 and do you know the rule and do you know the answer.
Answer by stanbon(75887)

80,77,71, blank,50,35,17 and do you know the rule and do you know the answer.
----
1st: 80
2nd: 80-3 = 77
3rd: 77- 2*3 = 71
4th: 71- 3*3 = 62
5th: 62- 4*3 = 50
6th: 50- 5*3 = 35
7th: 35- 6*3 = 17
---
Rule: A(n) = 80 - (n-1)3
OR
A(n) = 83-3n
=============
Cheers,
Stan H.
=============